### v3

# Dynamic Co-Optimization Compiler: Leveraging Multi-Agent Reinforcement Learning for Enhanced DNN Accelerator Performance

Abstract.
This paper introduces a novel Dynamic Co-Optimization Compiler (DCOC), which employs an adaptive Multi-Agent Reinforcement Learning (MARL) framework to enhance the efficiency of mapping machine learning (ML) models, particularly Deep Neural Networks (DNNs), onto diverse hardware platforms. DCOC incorporates three specialized actor-critic agents within MARL, each dedicated to different optimization facets: one for hardware and two for software. This cooperative strategy results in an integrated hardware/software co-optimization approach, improving the precision and speed of DNN deployments. By focusing on high-confidence configurations, DCOC effectively reduces the search space, achieving remarkable performance over existing methods. Our results demonstrate that DCOC enhances throughput by up to 37.95% while reducing optimization time by up to 42.2% across various DNN models, outperforming current state-of-the-art frameworks.
Multi-agaent reinforcement learning, Co-optimization, Hardware accelerators, Sampling, Throughput
journalyear: 2025copyright: rightsretainedconference: 30th Asia and South Pacific Design Automation Conference; January 20–23, 2025; Tokyo, Japanbooktitle: 30th Asia and South Pacific Design Automation Conference (ASPDAC ’25), January 20–23, 2025, Tokyo, Japandoi: 10.1145/3658617.3697547isbn: 979-8-4007-0635-6/25/01ccs: Computing methodologies Multi-agent reinforcement learningccs: Hardware Hardware accelerators
1. Introduction
The computational demands and complexity of deep neural networks (DNNs) have surged with their expanding applications across various industries such as autonomous driving, medical imaging, and natural language processing (Hossain et al., 2023; Ashrafi et al., 2024; Mokhtari Dowlatabad et al., 2023; Farsi et al., 2024; Razmara et al., 2024). Traditionally, performance optimization in frameworks like TensorFlow (Abadi et al., 2016) and PyTorch (Paszke et al., 2019) has relied on hand-optimized kernels like NVIDIA’s cuDNN and Intel’s MKL. However, these kernels struggle to adapt to the evolving demands of modern neural architectures (Ahn et al., 2020).
To address this, automated compilation frameworks like TVM (Chen et al., 2018), TensorComprehensions (Vasilache et al., 2018), and AutoTVM (Chen et al., 2018) have emerged, shifting from static, hand-tuned optimization to dynamic, algorithm-driven approaches. For instance, AutoTVM uses boosted trees (Chen and Guestrin, 2016) to efficiently explore the configuration space of neural network code. Despite these advances, optimizing complex models still requires substantial resources and compilation time, highlighting ongoing challenges in current optimization methodologies. Reinforcement learning (RL) has shown promise in exploring design search spaces and optimizing tasks (Zhang et al., 2022; Wang et al., 2022; Bakshi and Johnsson, 2023; Ahn et al., 2020), with multi-agent RL offering advantages like transfer learning and insights into interconnected decision-making systems (Du and Ding, 2021).
In this work, we propose an adaptive multi-agent reinforcement learning-based hardware/software co-optimization compiler called DCOC. Our compiler employs a multi-agent reinforcement learning (MARL) algorithm to concurrently optimize the mapping of the DNN model and the accelerator architecture. The use of MARL effectively manages the expansive search space of hardware and software configurations, leveraging the collaborative strengths of multiple agents (Yu et al., 2022), which is crucial due to the significant impact of hardware-software interplay on overall performance.
DCOC features three specialized agents: two focusing on software and DNN configurations, and one optimizing the accelerator architecture. This actor-critic MARL approach enables an integrated and holistic optimization process with granular control over optimization strategies. The framework efficiently narrows the vast search space by employing a soft threshold to select high-confidence configurations, concentrating on the most promising candidates, thus reducing computational overhead and exploration time.
Utilizing a Centralized Training with Decentralized Execution (CTDE) strategy, DCOC allows adaptive adjustments to the changing requirements of DNN workloads. We assess the efficacy of DCOC on a set of DNN models with various architectures, comparing it with leading state-of-the-art approaches. The results demonstrate substantial improvements in throughput, showcasing the potential of MARL-based co-optimization in advancing compilation technologies.
2. Background
2.1. CTDE in MARL
Multi-agent reinforcement learning (MARL) extends traditional reinforcement learning by introducing multiple agents within a shared environment, necessitating methods to coordinate strategies towards collective objectives. The Centralized Training with Decentralized Execution (CTDE) framework balances collective intelligence during training with agent autonomy during execution (Lyu et al., 2021).
In CTDE, agents have access to global state information during centralized training, allowing them to learn cohesive policies that consider all agents’ states and actions. This approach effectively addresses cooperative and competitive tasks. Proximal Policy Optimization (PPO) (Schulman et al., 2017) has been adapted to this framework, resulting in Multi-Agent PPO (MAPPO) (Yu et al., 2022).
The MAPPO learning process involves three key components:
  • Critic learning: Enhances the centralized value function:
    (1) $\phi_{k+1}=\arg\min_{\phi}\frac{1}{|D_{k}|T}\sum_{\tau\in D_{k}}\sum_{t=0}^{T}\left(V_{\phi}(\mathbf{o}_{t},\mathbf{s}_{t},\mathbf{u}_{t})-\hat{R}_{t}\right)^{2}$
  • Generalized Advantage Estimation (GAE): Gauges the quality of the current action relative to the baseline:
    (2) $A_{t}=\sum_{t=0}^{\infty}(\gamma\lambda)^{t}\delta^{V}_{t+1}$
  • Policy learning: Updates the policy using the estimated advantage:
    (3) $L(\mathbf{o},\mathbf{s},\mathbf{u},\mathbf{u}^{\prime},\theta_{k},\theta)=$ $\min\bigg{(}\frac{\pi_{\theta}(\mathbf{u}|\mathbf{o})}{\pi_{\theta_{k}}(\mathbf{u}|\mathbf{o})}A^{\pi_{\theta_{k}}}(\mathbf{o},\mathbf{s},\mathbf{u}^{\prime}),$
    $\text{clip}\left(\frac{\pi_{\theta}(\mathbf{u}|\mathbf{o})}{\pi_{\theta_{k}}(\mathbf{u}|\mathbf{o})},1-\epsilon,1+\epsilon\right)A^{\pi_{\theta_{k}}}(\mathbf{o},\mathbf{s},\mathbf{u}^{\prime})\bigg{)}$
Empirical studies show that MAPPO can rival and sometimes outperform competitive off-policy algorithms (Yu et al., 2022), leading us to adopt MAPPO as our optimization algorithm in MARL.
2.2. Workflow for DNN Compilers
Efficiently compiling DNN models into machine code is crucial for performance across various hardware platforms. Compiler workflows optimize the DNN model in several stages (Allen and Kennedy, 2002). Initially, the compiler’s frontend performs general optimizations without considering specific hardware. Subsequently, backend optimizations tailor the code based on the target hardware’s characteristics.
Recent advancements like AutoTVM have introduced a fine-tuning step that uses hardware performance feedback to explore various configuration settings (”knobs”) and determine the most effective code configuration. The search for the optimal configuration ($\Theta^{*}$) is expressed as:
$\Theta^{*}=\underset{\Theta\in D}{\text{argmax}}\left\{f\left[\tau(\Theta)\right]\right\},$
where $f$ represents the performance metric, $\Theta$ denotes a specific set of knob settings, and $D$ is the allowed configuration space. The compiler’s goal is to discover the settings that maximize performance on the target hardware.
2.3. Related Work
In the domain of DNN compilers, adaptive and efficient auto-tuning mechanisms are essential for optimizing performance on varying hardware architectures (Rieber et al., 2022; Sadeghi et al., 2024; Ryu et al., 2022; Lu et al., 2022; Fayyazi et al., 2024; Dhakal et al., 2022; Azizi et al., 2024). AutoTVM (Chen et al., 2018) utilizes a machine learning-based cost model, specifically XGBoost (Chen and Guestrin, 2016), but involves a large search space and substantial measurement overhead.
MetaTune (Ryu and Sung, 2021) addresses these limitations by leveraging meta-learning to quickly adapt to new optimization spaces, demonstrating improved inference times. Glimpse (Ahn et al., 2022) integrates mathematical embeddings of hardware specifications, guiding the search algorithm towards higher performance subspaces using Bayesian optimization.
CHAMELEON (Ahn et al., 2020) employs reinforcement learning to minimize invalid configurations and costly hardware measurements, reducing optimization time and improving inference times compared to AutoTVM. NaaS (Zhou et al., 2022) jointly optimizes neural network architectures and hardware accelerators, parameterizing both in a unified search space using PyGlove. PRIME (Kumar et al., 2021) introduces a data-driven method for designing hardware accelerators, using logged simulation data to construct a robust surrogate model, significantly reducing simulation time.
Comparing these methods, AutoTVM lays the groundwork for machine learning-based auto-tuning. MetaTune and Glimpse advance this with meta-learning and hardware-aware strategies, respectively. CHAMELEON enhances the reinforcement learning approach with improved sampling methods. NaaS and PRIME introduce accelerator co-design and offline data-driven optimization but lack reinforcement learning, making them slower in compilation times.
Our approach differs from existing methods like CHAMELEON, Glimpse, and MetaTune—which reduce compilation times through various optimization techniques—and co-design methods like NaaS and PRIME—which enhance system throughput. Our analysis highlights the advantages of MARL in efficiently navigating the design space in modern DNN applications, capturing hardware and software interdependencies to achieve optimal performance and throughput in real-world applications.
3. Proposed DCOC
Existing systems often fail to optimize both hardware and software components of DNNs simultaneously, focusing primarily on software parameters, while co-design approaches are too time-consuming to compile. To address these limitations, we identify two key challenges:
  1. (1)
    Enhancing the search algorithm to include hardware architecture and parameters as adjustable factors.
  2. (2)
    Refining the sampling process to better capture the solution distribution and minimize non-feasible configurations.
We propose two main advancements in the co-optimizing compiler for DNNs. First, we integrate a multi-agent reinforcement learning (MARL) strategy into the search algorithm, enabling simultaneous optimization of hardware and software parameters. This approach leverages collective intelligence for a more holistic and efficient optimization. Second, we introduce a novel Confidence Sampling (CS) method to replace existing uniform (Chen et al., 2018) and adaptive sampling methods (Ahn et al., 2020), identifying configurations with a higher probability of success based on learned patterns.
Our primary optimization goal is to maximize throughput. The search space comprises configurations defined by various knobs, including tile dimensions for batch size, input channels, output channels, and threading options. By systematically exploring these configurations, our framework can adapt to different hardware platforms and optimize performance effectively.
3.1. Overview of DCOC
Figure 1. Overall search flow of DCOC.
Figure  1 depicts the architecture of our sophisticated co-optimizing compiler, termed DCOC, which elucidates the process of optimization compilation. DCOC initiates with a code template, $\tau$, for each layer of the neural network and a corresponding design space $D_{\Theta}$, delineating the range of possible configurations. The compiler then engages in an optimization process, navigating through the configuration space $\Theta$ to ascertain the optimal code configuration, represented as $\tau(\Theta^{*})$.
DCOC maneuvers through the design space, employing a cost model as a surrogate for direct hardware measurements. The MARL Exploration module generates an initial set of candidate configurations $S_{\Theta}$. Through the CS method, DCOC refines $S_{\Theta}$ to produce a more focused subset $S^{\prime}_{\Theta}$, which contains a condensed yet highly promising set of configurations. The configurations in $S^{\prime}_{\Theta}$ are then passed to the MARL Code Generator module, which incorporates them with the input template $\tau$ to create a series of potential executable codes $\tau(\Theta)$. These are subsequently deployed on the hardware for empirical runtime evaluations. The hardware runtimes provide a measure of the configurations’ fitness, captured by a fitness function $f$, which informs the update of the cost model and enhances the exploration in subsequent iterations. After several iterations, the process converges to some $\tau(\Theta^{*})$ that achieves optimal fitness, characterized by the shortest hardware execution runtime. This configuration $(\Theta^{*})$ is then selected as the output for that network layer.
3.2. MARL Exploration
Figure 2. High-level view of MARL Exploration Module. Each Agent has a policy network and, based on the centralized critic feedback, it will do an action in its own environment.
In our work, we implement a MARL Exploration module (As shown in Figure 2) that employs a Multi-Agent Reinforcement Learning (MARL) strategy designed to optimize the configurations of DNN architectures and hardware simultaneously. This module utilizes the principle of Centralized Training and Decentralized Execution (CTDE), allowing agents to learn collaboratively while operating independently during the execution phase to maximize the fitness function, $f$, of configurations within the configuration space, $S_{\Theta}$. MARL’s ability to manage complex decision-making environments makes it particularly suitable for navigating the high-dimensional configuration space of DNNs.
The MARL Exploration strategy employs three agents, each equipped with a policy network, as detailed in Table 1. Additionally, there is a centralized critic network (value network) that, along with the policy networks, is implemented as a Multi-Layer Perceptron (MLP). The policy network directs each agent to propose adjustments to the configuration knobs (as shown in Table 2) within its assigned portion of the design space. In contrast, the value network estimates the potential value of these adjustments. A shared central critic aids in evaluating the global state, facilitating collective learning during the training phase, while each agent makes decisions independently during execution, adhering to the CTDE paradigm.
Table 1. Roles and Responsibilities of Different Agents
Agent Type Description
Scheduling
Agent
Focuses on task parallelization and distribution for effective scheduling.
Mapping
Agent
Divides and processes data dimensions (tensor height and width) for optimal hardware computation mapping.
Hardware
Agent
Adjusts parameters for dividing and processing tensor components (batches, input and output channels) to optimize hardware utilization based on resource counts and performance levels.
The optimization unfolds through several episodes, each consisting of multiple search steps. During these steps, the agents evaluate the current configuration and independently decide on actions to improve subsequent configuration fitness, guided by shared and individual observations. The process detailed in Algorithm 1 outlines how each agent interacts within the CTDE, iterating through configurations to enhance the network performance.
Algorithm 1 CTDE MARL Exploration for DNN Configuration Optimization
1:Initialize centralized critic and individual policy networks for each agent
2:for each episode do
3:     Initialize a set of configurations, $S_{\Theta}$
4:     for each search step in the episode do
5:         for each agent do
6:               Observe the current configuration, $\Theta$, and receive shared insights from the centralized critic
7:               Independently chooses an action based on the policy network and local observations
8:              Apply the action to update the configuration
9:              Estimate the new configuration’s local value
10:         end for
11:          Collectively evaluate all new configurations using the cost model
12:          Update the centralized critic based on global performance feedback
13:          Individually update each agent’s policy network based on local and shared feedback
14:     end for
15:      Determine the configuration with the highest fitness from collective insights
16:end for
17:Return the optimal configuration, $\tau(\Theta^{*})$
Each episode integrates the evaluations from a cost model that serves as a surrogate for direct hardware performance measurements. This model’s feedback is used to update the centralized critic, enhance the shared knowledge base, and refine the individual policy networks, refining each agent’s decision-making capabilities based on global and local feedback. As the episodes progress, the MARL Exploration Module becomes increasingly adept at identifying configurations that yield the best performance, ultimately converging on $\tau(\Theta^{*})$ with the optimal fitness, $f$.
Applying this CTDE framework within the MARL context, coupled with the PPO optimization algorithm, represents a significant advancement in the auto-tuning of DNN and hardware configurations. This approach systematically improves the configuration search process, significantly enhancing DNN performance while reducing the computational overhead.
3.2.1. Cost Model and Central Critic Update
In the DCOC framework, the optimization goal is to improve throughput, with the cost model reflecting the inverse of execution time. The central critic, a key component of our MARL setup, updates its policy based on the aggregated feedback from all agents, continuously optimizing the global state evaluation. The critic’s learning process is guided by the update rule, which is mathematically expressed in section 2.1, which minimizes the mean squared error between the predicted values and the actual rewards obtained.
3.2.2. Incorporating Hardware and Software Constraints
To integrate constraints related to hardware, such as area limitations, or software, such as memory usage, into the MARL framework, a penalty term can be incorporated into the reward function. This approach adjusts the reward based on the degree to which the constraints are violated, effectively guiding the agents to prefer configurations that adhere to these constraints. For instance, a penalty function $P$ can be defined as follows:
$P(\Theta)=$ $\lambda\left(\max(0,\text{area}(\Theta)-\text{area}_{\text{max}})\right.$
(4) $\left.+\max(0,\text{memory}(\Theta)-\text{memory}_{\text{max}})\right)$
where $\lambda$ is a scaling factor that adjusts the impact of the penalty, $\text{area}(\Theta)$ and $\text{memory}(\Theta)$ are the area and memory usage of the configuration $\Theta$, and $\text{area}_{\text{max}}$ and $\text{memory}_{\text{max}}$ are the respective maximums.
The reward function, modified to include this penalty, becomes:
(5) $R_{t}=\frac{1}{\text{execution time}(\Theta)}-P(\Theta)$
This modification ensures that the MARL agents not only seek to optimize performance but also adhere to specified design constraints, balancing efficiency with practical deployment considerations. In this work, without loss of generality, we considered VTA++ (Banerjee et al., 2021) as the HW platform. The VTA++ architecture is comparable to other hardware platforms, such as GPUs and FPGAs, due to its configurable nature and similarity in processing elements (Banerjee et al., 2021). This similarity ensures that the optimization strategies and results obtained from VTA++ can be generalized and applied to these other hardware platforms, validating the broad applicability of our framework.
Table 2. Knobs in the design space to optimize convolution layers (These knobs make a search space as big as O($2^{12}$))
Agent Type Knobs
Hardware Design Agent
(Hardware Optimizer)
- Tile across batch size (tile_b)
- Tile across input channels (tile_ci)
- Tile across output channels (tile_co)
Scheduling Agent
(Software Optimizer)
- Horizontal threading (h_threading)
- Output channel threading (oc_threading)
Mapping Agent
(Software Optimizer)
- Tile across height (tile_h)
- Tile across width (tile_w)
3.3. Confidence Sampling
In DNN optimization domain, the ability to efficiently sample configuration spaces without costly hardware measurements is essential for rapid compilation. We introduce the Confidence Sampling (CS) method. This technique is designed to leverage the probability distribution of configurations’ fitness judiciously to select a subset likely to yield high-performing network configurations.
Although the Confidence Sampling method operates on the principles of probability-guided selection and shares some conceptual similarities with Importance Sampling (IS) (Tokdar and Kass, 2010) in statistical methods, it is distinct in its purpose and application. In IS, samples are re-weighted based on their probability density to approximate expected values more efficiently. Confidence Sampling, however, specifically focuses on reducing the number of hardware measurements required when exploring the design space. By evaluating the configurations’ values, this method synthesizes a subset that not only encapsulates the diversity of the configuration space but also emphasizes the high-confidence regions, i.e., areas where the likelihood of encountering an (near-) optimal configuration is high.
The core of the CS method lies in its algorithmic approach to filtering and enhancing the set of configurations for subsequent optimization iterations. To formalize this process, we introduce Algorithm 2, which captures the essence of the CS method within the optimization pipeline. The process involves the following steps:
  1. (1)
    Evaluate Configurations: Each configuration’s value is estimated using the output of the value network (critic network) to construct a probability distribution over the configuration space. (Line 2)
  2. (2)
    Probability-Guided Selection: Configurations are sampled based on this distribution, prioritizing those with higher estimated values. (Line 3-4)
  3. (3)
    Confidence Assessment: A dynamic thresholding mechanism is applied to determine the confidence level of each selected configuration. (Line 5)
  4. (4)
    Synthesis of Configurations: Lower-confidence configurations are replaced with synthesized ones by combining each parameter’s most frequently occurring settings across the sampled configurations. (Line 6-7)
Algorithm 2 Confidence Sampling for DNN Optimization
1:procedure ConfidenceSampling($\mathcal{S}_{\Theta}$, $value\_network$, $N_{select}$)
2:      $\mathcal{V}_{preds}\leftarrow value\_network.predict(\mathcal{S}_{\Theta})$triangleright Estimate values for all configurations
3:      $\mathcal{P}\leftarrow softmax(\mathcal{V}_{preds})$triangleright Convert values to a probability distribution
4:     $\mathcal{C}_{selected}\leftarrow\textsc{SelectConfigurations}(\mathcal{P},N_{configs})$
5:     $threshold\leftarrow\textsc{ComputeDynamicThreshold}(\mathcal{V}_{preds})$triangleright Compute an adaptive threshold
6:     $\mathcal{C}_{high\_conf}\leftarrow\{c\in\mathcal{C}_{selected}\mid\mathcal{V}_{preds}[c]>threshold\}$
7:     return $\mathcal{C}_{high\_conf}$
8:end procedure
9:function SelectConfigurations($\mathcal{P}$, $N_{configs}$)
10:      $indices\leftarrow$ sample indices from $[1,\ldots,|\mathcal{P}|]$ with probability $\mathcal{P}$
11:     return $\mathcal{S}_{\Theta}[indices[1:N_{configs}]]$
12:end function
13:function ComputeDynamicThreshold($\mathcal{V}_{preds}$)
14:     return $median(\mathcal{V}_{preds})$
15:end function
By applying the CS method, we have observed a reduced number of required configurations, as shown in Figure  3. Moreover, we can observe the sampling process gravitates towards configurations that demonstrate superior performance over time, highlighting the efficacy of this method.
Figure 3. Configurations over time for ResNet-18 model a) before and b) after applying the CS method.
4. Experimental Results
4.1. Experimental Setup
We have incorporated DCOC within the TVM framework to thoroughly evaluate its components, contrasting its performance with that of AutoTVM and CHAMELEON. This integration allows for a comprehensive, end-to-end assessment of DCOC across a suite of advanced deep-learning models, including AlexNet, VGG-11, VGG-13, VGG-16, VGG-19, ResNet-18, and ResNet-34. All models were benchmarked using the same AutoTVM compilation duration to ensure a rigorous and equitable comparison. Further, we have employed the source code of the CHAMELEON (Ahn et al., 2020) to compare its efficacy with AutoTVM and DCOC. As mentioned above, AutoTVM and CHAMELEON do not support hardware configuration exploration. Hence, we also added the hardware architecture agent for them to have a fair comparison. Performance tests were conducted on a robust platform powered by a 3.4 GHz AMD EPYC 7763 64-Core Processor, providing a stable and powerful environment for these evaluations. As discussed before, we used the VTA++ simulator as our target hardware.
In our setup, each agent is equipped with its own policy network, and a centralized value network is the critic for all agents. Both networks are implemented as Multi-Layer Perceptrons (MLPs) using TensorFlow (Abadi et al., 2016) that their details are:
  • Policy Network: Each agent’s policy network consists of a single hidden layer with 20 neurons using ReLU activation functions. The output layer uses softmax activation to generate a probability distribution over actions, allowing for decision-making informed by learned policies.
  • Value Network: The centralized critic, or value network, employs three hidden layers with 20 neurons each, using tanh activation functions. This design helps stabilize learning by providing a continuous estimate of state values, which guides collective policy adjustments.
Hyper-parameter tuning is crucial in optimizing the performance of machine learning tools and models (Yu and Zhu, 2020). To facilitate transparency and reproducibility, we have detailed the hyper-parameters utilized in our evaluation in Table 3. These hyper-parameters were meticulously tuned offline to optimize model performance, and they are the same as those used in CHAMELEON (Ahn et al., 2020). For the hyper-parameters listed in Table 4, we adhered to the values specified in the AutoTVM study to maintain consistency and ensure a fair comparison. Similarly, we adopted the hyper-parameter settings from the MAPPO paper (Yu et al., 2022) for the MARL exploration module, aligning our methodologies with established research to validate our findings effectively.
Table 3. Hyper-parameters (HPs) used in DCOC.
HP Value Description
iteration_opt 16
Total iterations for the optimization cycle
(equivalent to 1000 hardware measurements)
modeGBT xgb-reg Loss function type utilized in the cost model
bGBT 64 Planning’s maximum batch size in GBT (Chen et al., 2018)
episode_rl 128 Number of episodes for the RL process
step_rl 500 Maximum # of steps in a single RL episode
Table 4. Hyper-parameters (HPs) used in AutoTVM
HP Value Description
$\Sigma(b_{GBT})$ 1000 Total count of hardware measurements
$mode_{GBT}$ xgb-reg Loss function type for the cost model
$b_{GBT}$ 64 Batch size for planning in GBT (Chen et al., 2018)
$n_{sa}$ 128 Count of Markov chains in parallel SA
$step_{sa}$ 500 Maximum # of steps in a single SA run
  • SA: Simulated Annealing.
4.2. End-to-end Evaluation
Figure 4 Compares the achieved throughput of different frameworks over AutoTVM. As the results illustrate, the significant reduction in mean inference time (advancement in throughput) achieved by DCOC compared to AutoTVM and CHAMELEON for various DNN models: AlexNet, VGG-11, VGG-13, VGG-16, VGG-19, and ResNet-18, and ResNet-34. On average, our methodology demonstrates a 1.17× improvement in throughput. Table 5 reports detailed numerical results supporting these findings.
Figure 4. Comparing the achieved throughput of different frameworks over AutoTVM on VTA++.
Table 5. Mean inference times for different frameworks on VTA++ (in seconds)
Model AutoTVM CHAMELEON DCOC
ResNet-18 1.730 611.730611.730\,611.730 61 1.595 951.595951.595\,951.595 95 1.25448
ResNet-34 3.634 093.634093.634\,093.634 09 3.537 803.537803.537\,803.537 80 2.66561
AlexNet 0.541 160.541160.541\,160.541 16 0.589 810.589810.589\,810.589 81 0.42160
VGG11 4.952 294.952294.952\,294.952 29 4.212 834.212834.212\,834.212 83 4.20685
VGG13 6.192 896.192896.192\,896.192 89 6.163 506.163506.163\,506.163 50 6.03418
VGG16 8.558 788.558788.558\,788.558 78 8.730 898.730898.730\,898.730 89 8.34791
VGG19 11.291 5011.2915011.291\,5011.291 50 11.288 3911.2883911.288\,3911.288 39 11.27800
Figure 5. Comparing the compilation time of different frameworks (The percentages show the speedup of DCOC compared to AutoTVM).
While achieving improvement in throughput, DCOC also needs less compilation time compared to AutoTVM; Figure 5 shows that DCOC speeds up the optimization time (the entire process of compiling) up to 42.2%. Figure 6 shows DCOC’s comparative output code performance relative to other frameworks. Notably, DCOC achieves convergence to the peak GFLOPS value, equivalent to those reached by AutoTVM and CHAMELEON (without adaptive sampling), but with greater efficiency. This is accomplished with fewer hardware measurements and at a faster rate, underscoring the efficacy of the CS method.
Figure 6. Comparing the compiled code performance (GFLOPS per Trial) of different frameworks for ResNet-18 model.
5. Conclusion
This paper introduced DCOC, an innovative multi-agent reinforcement learning-based compiler that significantly advances the field of DNN accelerator design. By integrating a multi-agent system with specialized roles—two agents focused on software optimization and one on hardware—DCOC addressed the intricacies of DNN architecture and effectively navigated the vast configuration space to enhance performance and efficiency. Employing the MARL Exploration module and Confidence Sampling method, DCOC reduced the computational overhead and expedited the optimization process while increasing the throughput.